# On the Theory and Practice of GRPO: A Trajectory-Corrected Approach with Fast Convergence

### v1

@@ 4 Trajectory-Level Importance-Corrected GRPO (TIC-GRPO)
$\mathcal{L}_{\text{TIC-GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ $=\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\min\left\{{\color[rgb]{0,0,1}w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})}A_{i},\overline{\operatorname{clip}}\bigl{(}{\color[rgb]{0,0,1}w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})},\epsilon_{0}\bigr{)}A_{i}\right\}$
$\quad-\beta\,\mathrm{KL}\bigl{(}\pi_{\theta}\|\,\pi_{\theta_{\text{ref}}}\bigr{)}.$ (7)
We highlight the only modified component $w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})$ in blue; all other parameters remain identical to those in GRPO and follow the same definitions provided in Section 2.2. Similarly, we present the update rule under a fixed old policy $\pi_{\text{old}}:$
$\theta_{n+1}=\theta_{n}+\eta_{n}\widehat{\nabla}\mathcal{L}_{\text{TIC-GRPO}}(\theta_{n},\theta_{\text{old}},\theta_{\text{ref}}),$
@@
$\nabla\mathcal{L}_{\text{TIC-GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$ $=\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\nabla\left(\min\left\{w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})A_{i},\overline{\operatorname{clip}}\bigl{(}w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}}),\epsilon_{0}\bigr{)}A_{i}\right\}\right)}_{\emph{Gradient Term}}$
$\quad\underbrace{-\beta\nabla\left(\mathrm{KL}\bigl{(}\pi_{\theta}\|\,\pi_{\theta_{\text{ref}}}\bigr{)}\right)}_{\emph{Penalty Term}}.$ (8)
@@
Input: reward function $r(\cdot)$; initial parameters $\theta$; reference parameters $\theta_{\mathrm{ref}}$; group size $|G|;$ clipping threshold $\epsilon_{0}$; KL coefficient $\beta$;
inner iterations $K$; learning-rate schedule $\{\eta_{n}\}_{1:N}$
@@ 10 $A_{i}\leftarrow\dfrac{r_{i}-\mu_{G}}{\sigma_{G}+\delta}$;
14    $\mathcal{L}\leftarrow\frac{1}{|G|}\sum_{i,j}\left(\min\left\{w_{i,j}A_{i},\,\overline{\operatorname{clip}}\bigl{(}w_{i,j},\,\epsilon_{\text{low}},\,\epsilon_{\text{high}}\bigr{)}A_{i}\right\}\;+\beta\,D_{\mathrm{KL}}\!\Bigl{(}\pi_{\theta}\bigl{(}\cdot\mid s^{(i)}_{j-1}\bigr{)}\,\|\,\pi_{\theta_{\mathrm{ref}}}\bigl{(}\cdot\mid s^{(i)}_{j-1}\bigr{)}\Bigr{)}\right)$;
@@
16    $\theta\leftarrow\theta+\eta_{n}\,\hat{\nabla}\mathcal{L}$;
@@
$\mathcal{B}^{+}(s_{T}^{(i)},\theta,\theta_{\text{old}}):=\left\{w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})\leq 1+\epsilon_{0},\ A_{i}\geq 0\right\},$
$\mathcal{B}^{-}(s_{T}^{(i)},\theta,\theta_{\text{old}}):=\left\{w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})\geq 1-\epsilon_{0},\ A_{i}\leq 0\right\},$
along with their union event
@@
$\quad\text{\emph{Gradient Term}}\mathop{=}^{(*)}\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}(s_{T}^{(i)},\theta,\theta_{\text{old}})}w’(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot A_{i}$
$=\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}w’(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot A_{i}$
$\quad-\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}^{c}(s_{T}^{(i)},\theta,\theta_{\text{old}})}w’(s_{t}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot A_{i}$
${=}\frac{1}{\overline{\sigma}_{G}}\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}w’(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(r_{i}-\overline{\mu}_{G}\right)}_{\widetilde{\nabla}J(\theta)}$
$\quad+\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}w’(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(A_{i}-\frac{r_{i}-\overline{\mu}_{G}}{\overline{\sigma}_{G}}\right)}_{\text{\emph{Sampling Error}}}$
$\quad\underbrace{-\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{B}^{c}(s_{T}^{(i)},\theta,\theta_{\text{old}})}w’(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot A_{i}}_{\text{\emph{Clip Error}}}.$ (9)
@@
$\operatorname{\mathbb{E}}_{G}\left[\widetilde{\nabla}J(\theta_{\text{old}})\right]$ $=\operatorname{\mathbb{E}}_{G}\left[\frac{1}{|G|}\sum_{i=1}^{|G|}w’(s_{T}^{(i)},\theta,\theta_{\text{old}})\left(\nabla\log\left(\prod_{t=1}^{T}\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\right)\right)\cdot(r_{i}-\overline{\mu}_{G})\right]$
$=\operatorname{\mathbb{E}}_{s_{T}\sim\pi_{\theta_{\text{old}}}}\left[\frac{\operatorname{\mathbb{P}}_{\theta}(s_{T}|s_{0})}{\operatorname{\mathbb{P}}_{\theta_{\text{old}}}(s_{T}|s_{0})}\left(\nabla\log\operatorname{\mathbb{P}}_{\theta}(s_{T}|s_{0})\right)r(s_{T})\right]-0$
@@ 5 Convergence Results
$\theta_{n,s+1}=\theta_{n,s}+\eta_{n}\hat{\nabla}\mathcal{L}_{\text{GRPO}}(\theta_{n,s},\theta_{n,0},\theta_{\text{ref}}),\quad(s=0,1,\dots,K).$ (10)
@@
$\theta_{n,s+1}=\theta_{n,s}+\eta_{n}\hat{\nabla}\mathcal{L}_{\text{TIC-GRPO}}(\theta_{n,s},\theta_{n,0},\theta_{\text{ref}}),\quad(s=0,1,\dots,K).$ (11)

### v2

@@ 4 Trajectory-Level Importance-Corrected GRPO (TIC-GRPO)
Moreover, we employ a minor technical modification to the standard clipping mechanism used in importance sampling. The original clipping strategy, as defined in Eq. 2.2, i.e.,
$\min\left\{{w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})}A_{i},\operatorname{clip}\bigl{(}{w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})},\epsilon_{\text{low}},\epsilon_{\text{high}}\bigr{)}A_{i}\right\},$
treats the sign of the estimated advantage $A_{i}$ separately: when $A_{i}\geq 0$, the importance weight is clipped from above by $1+\epsilon_{\mathrm{high}}$, whereas when $A_{i}<0$, it is clipped from below by $1-\epsilon_{\mathrm{low}}$. However, we observe that retaining only the lower bound $1-\epsilon_{\mathrm{low}}$ while leaving the upper bound unconstrained fails to reduce the variance of the policy gradient estimator effectively even when $A_{i}<0.$ Motivated by this, we adopt a modified clipping scheme in which only the upper bound is enforced, as follows:
$\min\left\{{w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})},\operatorname{clip}\bigl{(}{w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})},\epsilon_{\text{low}},\epsilon_{\text{high}}\bigr{)}\right\}A_{i}.$
It is evident that under our modified clipping scheme, the importance sampling ratio is truncated from above at $1+\epsilon_{\mathrm{high}}$, regardless of the sign of the estimated advantage $A_{i}$. In particular, ratios exceeding $1+\epsilon_{\mathrm{high}}$ are clipped, while the lower bound $1-\epsilon_{\mathrm{low}}$ becomes inactive and thus plays no role in the computation. This modification leads to a more effective reduction in the variance of the policy gradient estimator, which in turn results in improved empirical performance.
For notational convenience, we denote:
$\overline{\text{clip}}(x,\epsilon_{\text{high}}):=\min\left\{x,\operatorname{clip}\bigl{(}x,\epsilon_{\text{low}},\epsilon_{\text{high}}\bigr{)}\right\}.$
@@
$\quad\mathcal{L}_{\text{TIC-GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$
$=\frac{1}{|G|}\sum_{i=1}^{|G|}\frac{1}{|s_{T}^{(i)}|}\sum_{t=1}^{T}\overline{\text{clip}}(w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}}),\epsilon_{\text{high}})A_{i}-\beta\text{D}_{\text{KL}}\bigl{(}\pi_{\theta}\|\pi_{\theta_{\text{ref}}}\bigr{)}.$ (9)
All other parameters remain identical to those in GRPO and follow the same definitions provided in Section 2.2. Similarly, we present the update rule under a fixed old policy $\pi_{\text{old}}:$
$\theta_{n+1}=\theta_{n}+\eta\widehat{\nabla}\mathcal{L}_{\text{TIC-GRPO}}(\theta_{n},\theta_{\text{old}},\theta_{\text{ref}}),$
@@
$\quad\nabla\mathcal{L}_{\text{TIC-GRPO}}(\theta,\theta_{\text{old}},\theta_{\text{ref}})$
$=\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\frac{1}{|s_{T}^{(i)}|}\sum_{t=1}^{T}\nabla\left(\overline{\operatorname{clip}}(w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}}),\epsilon_{\text{high}})\right)A_{i}}_{\emph{Gradient Term}}-\underbrace{\beta\nabla\left(\text{D}_{\text{KL}}\bigl{(}\pi_{\theta}\|\pi_{\theta_{\text{ref}}}\bigr{)}\right)}_{\emph{Penalty Term}}.$ (10)
@@
Input: reward function $r(\cdot)$; initial parameters $\theta$; reference parameters $\theta_{\mathrm{ref}}$; group size $|G|;$ clipping threshold $\epsilon_{\text{low}},\epsilon_{\text{high}}$; KL coefficient $\beta$;
inner iterations $K$; learning rate $\eta$
@@ 10 $A_{i}\leftarrow\dfrac{r_{i}-\mu_{G}}{\sigma_{G}+\delta}$;
14    $\mathcal{L}\leftarrow\frac{1}{|G|}\sum_{i=1}^{|G|}\frac{1}{|s_{T}^{(i)}|}\sum_{j=1}^{T}\left(\overline{\text{clip}}(w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}}),\epsilon_{\text{high}})A_{i}-\beta\text{D}_{\mathrm{KL}}\Bigl{(}\pi_{\theta}\bigl{(}s^{(i)}_{j}\mid s^{(i)}_{j-1}\bigr{)}\|\pi_{\theta_{\mathrm{ref}}}\bigl{(}s^{(i)}_{j}\mid s^{(i)}_{j-1}\bigr{)}\Bigr{)}\right)$;
@@
16    $\theta\leftarrow\theta+\eta\hat{\nabla}\mathcal{L}$;
@@
$\mathcal{D}(s_{T}^{(i)},\theta,\theta_{\text{old}}):=\left\{w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})\leq 1+\epsilon_{\text{high}}\right\},$
@@
$\quad\text{\emph{Gradient Term}}\mathop{=}^{(*)}\frac{1}{|G|}\sum_{i=1}^{|G|}\frac{1}{|s_{T}^{(i)}|}\sum_{t=1}^{T}\bm{1}_{\mathcal{D}(s_{T}^{(i)},\theta,\theta_{\text{old}})}w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot A_{i}$
${=}\frac{T_{\theta_{\text{old}}}^{(-1)}}{\overline{\sigma}_{G}}\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(r_{i}-\overline{\mu}_{G}\right)}_{\widetilde{\nabla}J(\theta)}$
$\quad+\underbrace{\frac{T_{\theta_{\text{old}}}^{(-1)}}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{D}(s_{t}^{(i)},\theta,\theta_{\text{old}})}w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot\left(A_{i}-\frac{r_{i}-\overline{\mu}_{G}}{\overline{\sigma}_{G}}\right)}_{Sampling\ Error_{1}}$
$+\underbrace{\frac{1}{|G|}\sum_{i=1}^{|G|}\left(T_{\theta_{\text{old}}}^{(-1)}-\frac{1}{|s_{T}^{(i)}|}\right)\sum_{t=1}^{T}\bm{1}_{\mathcal{D}(s_{T}^{(i)},\theta,\theta_{\text{old}})}w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot A_{i}}_{Sampling\ Error_{2}}$
$\quad\underbrace{-\frac{T_{\theta_{\text{old}}}^{(-1)}}{\overline{\sigma}_{G}}\frac{1}{|G|}\sum_{i=1}^{|G|}\sum_{t=1}^{T}\bm{1}_{\mathcal{D}^{c}(s_{T}^{(i)},\theta,\theta_{\text{old}})}w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})\nabla\log\pi_{\theta}(s_{t}^{(i)}|s_{t-1}^{(i)})\cdot(r_{i}-\overline{\mu}_{G})}_{\text{\emph{Clip Error}}}.$ (11)
@@
$\quad\operatorname{\mathbb{E}}_{s^{(1)},\ s^{(2)},...,\ s^{(|G|)}\sim\pi_{\theta_{\text{old}}}}\left[\widetilde{\nabla}J(\theta_{\text{old}})|\mathscr{F}_{\theta_{\text{old}}}\right]$
$=\operatorname{\mathbb{E}}_{s^{(1)},\ s^{(2)},...,\ s^{(|G|)}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{|G|}\sum_{i=1}^{|G|}w^{\prime}(s_{T}^{(i)},\theta,\theta_{\text{old}})\left(\nabla\log\left(\prod_{t=1}^{T}\pi_{\theta_{\text{old}}}(s_{t}^{(i)}|s_{t-1}^{(i)})\right)\right)\cdot(r_{i}-\overline{\mu}_{G})\Bigg{|}\mathscr{F}_{\theta_{\text{old}}}\right]$
$=\operatorname{\mathbb{E}}_{s_{T}\sim\pi_{\theta_{\text{old}}}}\left[\frac{\operatorname{\mathbb{P}}_{\theta}(s_{T}|s_{0})}{\operatorname{\mathbb{P}}_{\theta_{\text{old}}}(s_{T}|s_{0})}\left(\nabla\log\operatorname{\mathbb{P}}_{\theta}(s_{T}|s_{0})\right)r(s_{T})\Big{|}\mathscr{F}_{\theta_{\text{old}}}\right]-0$
@@ 5 Convergence Results
$\theta_{n,s+1}=\theta_{n,s}+\eta\hat{\nabla}\mathcal{L}_{\text{GRPO}}(\theta_{n,s},\theta_{n,0},\theta_{\text{ref}}),\quad(s=0,1,\dots,K).$ (12)
@@
$\theta_{n,s+1}=\theta_{n,s}+\eta\hat{\nabla}\mathcal{L}_{\text{TIC-GRPO}}(\theta_{n,s},\theta_{n,0},\theta_{\text{ref}}),\quad(s=0,1,\dots,K).$ (13)